OpenStudy (anonymous):

Write down the first three partial sums for the infinite series. I will write the series with the equation box.

OpenStudy (anonymous):

Give me a sec 2 put teh equation in

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} [(-1)^n-1/ [n(n+1)]\]

OpenStudy (anonymous):

CORRECTION. The numerator shud be (-1) to the (n-1) power

OpenStudy (anonymous):

and the denominator is same: n(n+1)

OpenStudy (phi):

\[ \sum_{n=1}^{\infty} \frac{(-1)^{n-1}}{ n(n+1)} \]

OpenStudy (phi):

for the first term, replace n with 1

OpenStudy (phi):

-1^(0) = 1 1(1+2) = 2

OpenStudy (phi):

so 1/2

OpenStudy (phi):

*1(1+1) = 2

OpenStudy (anonymous):

wat does the star mean

OpenStudy (phi):

typo

OpenStudy (anonymous):

the asnwer is 1/2 though not 2

OpenStudy (phi):

Yes, Look up a few boxes. I mean to type -1^(0) = 1 1(1+1) = 2 where the first line is the top of the fraction, and the second line is the denominator

OpenStudy (anonymous):

oh ok i see

OpenStudy (phi):

try the next term, with n=2

OpenStudy (anonymous):

so i just write S sub 2 = -1/6

OpenStudy (phi):

-1^(2-1)= -1^1= -1 2(2+1)=2*3= 6 so -1/6 Yes

OpenStudy (anonymous):

NVR MIND teh 1/6

OpenStudy (phi):

You are just interpreting short hand, because math people are lazy and don't want to write down all the terms...

OpenStudy (anonymous):

so S sub 3 is = 1/12

OpenStudy (anonymous):

but to find teh partial sum dont u have tofind a pattern and then write it as S sub n = sum equation

OpenStudy (phi):

Yes, we should be adding them up. Did I forget to do that? so 1/2 - 1/6 = 3/6 - 1/6 = 2/6 = 1/3 for the 2nd partial sum

OpenStudy (phi):

and the 3rd should be 1/3+1/12 = 4/12+1/12 = 5/12

OpenStudy (anonymous):

i see =) tahnks. can i ask u to help me with another question about convergence/divergence?

OpenStudy (anonymous):

i will post it

OpenStudy (phi):

ok

OpenStudy (anonymous):

shud i put it here or post it on teh side

OpenStudy (phi):

new post. Otherwise, these things get too long.

OpenStudy (anonymous):

ok just give me a sec